# Attention-based Reinforcement Learning for Combinatorial Optimization: Application to Job Shop Scheduling Problem

### 29 Jan 2024 v1

@@ 6 Conclusions
We developed an attention-based reinforcement learner for job shop scheduling problems (ARLS) with a modified transformer architecture. A key point was o enable the learner to scale with problem instance size. The learner was trained with synthetic problem instances of a relatively small size in the number of jobs and machines and scaled to new problems of various sizes with both synthetic and benchmark datasets. We have obtained results in general outperforming results in previous studies in the experiments and compared to widely adopted heuristic rules. The model architecture is new, but the transformer is low in complexity (few layers), and further work might improve results with larger transformer architectures.

### 18 Mar 2024 v2

@@ 6 Conclusions
We developed an attention-based reinforcement learner for job shop scheduling problems with a modified transformer architecture. A key point of our model is that it enables the learner to scale with problem instance size. The learner was trained with synthetic problem instances of relatively small size in the number of jobs and machines and scaled to new problems of various larger sizes with both synthetic and benchmark datasets. We have obtained generally outperforming results compared to recent studies and widely adopted heuristic rules. The model architecture is new, but the transformer is low in complexity (few layers), and further work might improve results with larger transformer architectures.